OpenStudy (madhatter):

Which ordered pair is a solution of the equation y = 6x + 6 for x = –2. (1 point) (–2, –6) (18, –2) (–2,-3/4 ) (–2, 24)

OpenStudy (anonymous):

lol back sos

OpenStudy (shamim):

-2, -6

OpenStudy (campbell_st):

just substitute and calculate well x = -2 eliminates the 2nd choice so look at 1, 3, 4 e.g. (-2, 24) does 24 = 6*(-2) + 6 24 = -12 + 6 Nope so its not the last solution.. leaves 1 and 3

OpenStudy (anonymous):

ohhhhhhhhhhh I get it now..... the answer is -2 -6 is easier when u do it like before

OpenStudy (madhatter):

so its not -2, -6

OpenStudy (anonymous):

no it is -2 -6 i think

OpenStudy (madhatter):

kk thanks